OpenStudy (anonymous):

Which expression is equivalent to the given expression? 12(x – 17)

OpenStudy (the_fizicx99):

12 * x 12 * (-17) multiply 12*x and 12 * (-17)

OpenStudy (anonymous):

what about like 5 more please

OpenStudy (anonymous):

what expression is equivalent to the given expression? 13x – 29x

OpenStudy (the_fizicx99):

Subtract like terms, 13 - 29 = -16 So, -16x

OpenStudy (solomonzelman):

\[a(b-c)=ab-ac\]\[a(b+c)=ab+ac\]\[(a^5)~^6=a^{5+6}=a^{11}\]\[a^{-b}=1/a^b\]\[7x-4x=(7-4)x=3x\]\[8d+4d=(8+4)d=12d\]\[a^7/a^3=a^{7-3}=a^4\]more tips?

OpenStudy (anonymous):

that was so confusing

OpenStudy (the_fizicx99):

Those are rules lol, some are exponent, others are distributive property etc. you'll learn about those later on.

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 2 – 8x + 3x when x = 5

OpenStudy (anonymous):

i reallyneed to pass this thing can u please help me

OpenStudy (the_fizicx99):

8(5) + 3(5) 40 + 15 - 2 = ?

OpenStudy (anonymous):

53??

OpenStudy (solomonzelman):

\[Evaluate~~~~~the~~~~~expression,~~~~~using~~~~~the~~~~~given~~~~~value\]\[of~~~~~the~~~~~variable.\]\[5+11x-6x~~~~~~~~~~~~~~~~~~~~~when~~~~~~~x=2\]\[5+11(2)-6(2)\]\[5+22-12\]\[15\]can you do your problem now?

OpenStudy (the_fizicx99):

Yup .. @SolomonZelman lol you're value is incorrect

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. z ÷ 4 – 8z + 2(z – 5) when z = 12

OpenStudy (the_fizicx99):

Ohh thats different nvm sorry >.>

OpenStudy (anonymous):

so it was 53

OpenStudy (the_fizicx99):

Plug in 12 for z in your expression @ala

OpenStudy (solomonzelman):

you tagged someone else, @tHe_FiZiCx99

OpenStudy (the_fizicx99):

12 / 4 - 8(12) + 2(12 - 5)

OpenStudy (the_fizicx99):

lol #Fail >.>

OpenStudy (anonymous):

so or 8z woule i multiply 8 and 12 or divide them

OpenStudy (anonymous):

for*

OpenStudy (anonymous):

??

OpenStudy (the_fizicx99):

Yes, plug in 12 in the equation.. 12 / 4 - 8(12) + 2(12 - 5)

OpenStudy (anonymous):

so its ether 79 or 74 right??

OpenStudy (the_fizicx99):

-79*

OpenStudy (anonymous):

Combine like terms. 7k – 13k

OpenStudy (the_fizicx99):

They're both the same.. just subtract 7 - 13 = ??

OpenStudy (anonymous):

6

OpenStudy (anonymous):

or -6

OpenStudy (the_fizicx99):

No... 7 MINUS 13 .. its negative. You have $7 and you have to pay 13.. your in debt. Also negative. <- Example..